1/1

Joseph Ferdinand Berger,

An den Mond

Verkinto de tiu ĉi Germana poemo estas Joseph Ferdinand Berger (∗1894-10-07 – †1976-03-28). Arg-782-1566 (2013-01-19 14:50:39)

http://www.poezio.net/version?poem-id=782&version-id=1566